OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. Write the solution as an ordered pair (a, b). a - b = -5 and a + 2b = 1 (3, 2) (3, -2) (-3, -2) (-3, 2) no solution

OpenStudy (saifoo.khan):

alright.

OpenStudy (anonymous):

-3,2

OpenStudy (anonymous):

answer

OpenStudy (saifoo.khan):

-3, 2

OpenStudy (anonymous):

-b+1-2b=-5 b=2 putting b in other eq a=1-4 a=-3 -3,2